### v1

# Training Machine Learning models at the Edge: A Survey

Abstract
Edge Computing (EC) has gained significant traction in recent years, promising enhanced efficiency by integrating Artificial Intelligence (AI) capabilities at the edge. While the focus has primarily been on the deployment and inference of Machine Learning (ML) models at the edge, the training aspect remains less explored. This survey delves into Edge Learning (EL), specifically the optimization of ML model training at the edge. The objective is to comprehensively explore diverse approaches and methodologies in EL, synthesize existing knowledge, identify challenges, and highlight future trends. Utilizing Scopus’ advanced search, relevant literature on EL was identified, revealing a concentration of research efforts in distributed learning methods, particularly Federated Learning (FL). This survey further provides a guideline for comparing techniques used to optimize ML for edge learning, along with an exploration of different frameworks, libraries, and simulation tools available for EL. In doing so, the paper contributes to a holistic understanding of the current landscape and future directions in the intersection of edge computing and machine learning, paving the way for informed comparisons between optimization methods and techniques designed for edge learning.
Index Terms:
Machine Learning; Edge Computing; Edge AI; Edge Learning; On-Device Training; Artificial Intelligence; IoT.
I Introduction
In the past years, the fields of Artificial Intelligence (AI) Machine Learning (ML) has witnessed significant growth and advancements, demonstrating remarkable success across various industrial applications [1]. The essence of ML has always been shaped by the interplay between data and algorithmic models. Large quantities of data being often required to effectively train ML models. And traditionally, the datasets were collected in Cloud Storage Buckets, Large Databases and Data Lakes, and processed in central cloud servers, to train various machine learning and deep learning models.
Conversely, The rapid proliferation of smart devices and sensors in recent years has led to an explosion of data generation at the edge of the network. With edge devices generating vast quantities of data outside the traditional cloud servers. The raising awareness of the privacy and security concerns that come with sending these data to the cloud, as well the desire to optimize the bandwidth consumption on the increasing number of edge devices and reduce the computational load on cloud servers, have precipitated a paradigm shift towards Edge Computing (EC). In this context, computational processes are decentralized and migrated to edge devices, setting the stage for a novel intersection between machine learning and edge computing.
This paradigm shift has spurred an increasing interest towards Edge Machine Learning (Edge ML). A union between Machine Learning and Edge Computing, where ML models are deployed at the edge, closer to end devices, for both inference and training purposes. The latter, known as Edge Learning (EL), involves the training of ML models directly at the edge. Traditionally, ML models have relied heavily on centralized cloud infrastructure for both training and deployment. However, this approach poses several challenges, including high latency, significant communication overheads, and concerns around data privacy and security. By moving the computation closer to the source of the data, EL offers a promising solution to address these issues while enabling real-time decision-making and reducing the load on cloud resources. Furthermore, edge learning creates opportunities for innovative applications of ML, such as privacy aware recommendation, smart technologies, and autonomous vehicles, spanning across multiple industries from healthcare and manufacturing to the agriculture and space industry.
However, training ML models at the edge presents unique challenges due to the limited computational capabilities and memory constraints of edge devices. Moreover, despite the vast amounts of data at the edge in general. The typical edge device generates insufficient data to facilitate the training of Machine Learning models from the ground up. To overcome these challenges and constraints, various techniques have been proposed such as Federated Learning (FL), Knowledge Distillation (KD), and Transfer Learning (TL), etc. These methods aim to optimize ML models to fit within the constraints of edge devices, thereby rendering them suitable for training in resource-constrained environments, scenarios with low data availability, or through collaborative training across multiple edge devices that leverage their collective data.
This survey paper aims to provide a general overview on Edge Learning (EL), where we look into its methodologies, requirements, applications, challenges, and open research directions. We explore the state-of-the-art techniques and approaches used to train and optimize ML models in edge devices, highlighting their advantages and making a broad comparison between them. Additionally, we investigate the diverse range of applications that benefit from Edge Learning, and the frameworks, libraries and simulation tools used to optimize and enable edge learning.
This survey is separated into eight different parts, excluding this Introduction and the Conclusions section at the end. First, Section II defines the terms edge learning and edge devices in more details, followed by the comparison of our survey with the existing ones highlighting our contributions in Section III, and the requirements and metrics for training ML models at the edge in Section IV. In Section (V), we explore the techniques that are used to enable, optimize and accelerate edge learning, and then make a detailed comparison between them in Section V-E. We explore how are the different types of ML techniques (Unsupervised learning, reinforcement learning, etc) are integrated with EL in Section VI, for the purpose of using these learning types in EL or to optimize or to enable the training of other models. In Section VII, we explore the use cases and current applications of edge learning. Then in Section VIII, we present different tools, frameworks and libraries used to train ML models at the edge (VIII-A) and create simulations and emulations for EL in Section VIII-B. Finally, in section IX, we highlight some open challenges, and try to predict future trends and research directions on edge learning.
II Edge computing and Edge learning
In this section, we will introduce the concepts of Edge Computing, Edge Machine Learning and Edge Learning.
II-A Edge Computing
Edge Computing (EC) is a new computing paradigm that aims to address the limitations of traditional Cloud Computing models in handling large scale data generated by the increasing number of smart devices connected to the Internet. It involves performing calculations at the edge of the network, closer to the user and the source of the data. EC emphasizes local, small-scale data storage and processing, providing benefits such as reduced bandwidth load, faster response speed, improved security, and enhanced privacy compared to traditional Cloud Computing models [2].
Edge Computing addresses several limitations of Cloud Computing, that stem from the frequent communications needed between end/edge devices and cloud server, in the standard Cloud Computing paradigm and the reliance of storing data centrally, which might compromise the privacy or security of sensible data.
  • Reduced latency: EC brings data processing closer to the source, reducing the time it takes for data to travel to a centralized cloud server, thereby reducing latency and improving response times [3].
  • Bandwidth optimization: EC reduces the need for transmitting large amounts of data to centralized cloud servers, resulting in reduced bandwidth load and reduced network congestion [2].
  • Improved data privacy: EC allows for local data processing, reducing the need to transmit sensitive data to centralized cloud servers, thereby minimizing the risk of data breaches and unauthorized access [2].
  • Operational resilience: EC enables applications to continue functioning even in disconnected or low-bandwidth environments, ensuring operational resilience and reducing dependency on a centralized cloud infrastructure [3].
The figure 1 shows the general Architecture of EC, inspired by the ones proposed in [4] and [5]. We define edge devices as both edge servers and end devices, as well as other types of devices that weren’t specifically mentioned in the diagram, such as routers and routing switches.
Figure 1: A typical architecture of Edge computing
II-B Edge Learning
A promising development in the field of EC involves the integration of AI and ML. The concept of Edge Machine Learning (Edge ML) strives to facilitate the training and deployment of ML models directly on edge devices, encompassing both edge learning and edge inference. EL is characterized by the direct training of ML models on edge devices, thereby diminishing the reliance on centralized cloud infrastructure for model training. On the other hand, Edge Inference pertains to the deployment of ML models and the facilitation of inference on resource-constrained edge devices, regardless of whether the models were trained on the edge or within a cloud-based environment [6].
Most of the work done on Edge ML today involves edge inference [7, 8], however Edge Learning (EL) is still a promising paradigm, as it allows for localized model training, tailored to the specific requirements and resource constraints of edge devices, making it ideal for application requiring privacy preservation, and model tailoring for specific use cases.
Edge Learning employs various strategies, most of them are either categorized as distributed or collaborative learning methods that aim to distribute the training of ML model to multiple edge devices, such as federated or split learning; and on-device learning that aim to train a ML model on a single edge device, using either optimization or fine-tuning techniques if needed.
In this survey, we will delve into both on-device learning and distributed learning on edge devices. Distributed learning in EL is defined as the training of ML models collaboratively across multiple edge devices. While on-device learning, refers to the training of ML models in a single edge device. To ensure clarity, our definition of edge devices also encompasses edge servers, network elements, and end devices. We comprehensively address machine learning model training across all these devices.
III Comparison with existing surveys
There has been considerable surveys and reviews about Edge ML and Edge AI that attempts to define the field and present the different approaches that exist for AI in Edge. Most of them focus on edge inference or focus on a single aspect of edge learning, such as Federated Learning [9, 10] or on-device training [11, 12].
As defined in Section I, the purpose of this survey is to give a complete overview on the training ML models on the edge. This survey has multiple contributions that we will use to compare with the existing ones:
  1. 1.
    Explore techniques: Exploring the different techniques used to train or facilitate the training of ML models on edge devices.
  2. 2.
    Metrics for Edge Learning: Defining metrics to evaluate and compare EL approaches, and also highlight the requirements necessary to use edge learning in real life scenarios.
  3. 3.
    Compare techniques: Compare the different learning techniques previously presented based on the metrics and requirements, as well as the usage and popularity of the techniques in academia in the context of EL.
  4. 4.
    Explore Types of ML: Explore the different types of ML such as unsupervised or reinforcement learning in the context of EL.
  5. 5.
    Explore tools and libraries: Explore the tools and libraries used for training ML models on edge devices and providing simulations and emulators for EL.
  6. 6.
    Use-cases and applications: Present different use-cases and applications for edge learning explored in Academia.
The Table I present the relevant studies related to Edge ML, and compare them to our survey based on the aforementioned points.
TABLE I: Summary of Edge Machine Learning related surveys
Survey Year Explore techniques Metrics for Edge Learning Compare techniques Explore Types of ML Explore tools and libraries Use-cases and applications
Chen et al. [4]
2019
Wang et al. [13]
2020
Shi et al. [14]
2020
Leon Veas et al. [15]
2021
Dhar et al. [12]
2021
$\circ$
$\circ$
Tak et al. [16]
2021
$\circ$
Murshed et al. [17]
2022
Abreha et al. [9]
2022
$\circ$
$\circ$
$\circ$
Boobalan et al. [10]
2022
$\circ$
$\circ$
Joshi et al. [18]
2022
Cai et al. [19]
2022
Cui et al. [20]
2022
$\circ$
$\circ$
Imteaj et al. [21]
2022
$\circ$
$\circ$
$\circ$
Singh et al. [22]
2023
Li et al. [23]
2023
Hua et al. [24]
2023
Zhu et al. [11]
2023
$\circ$
$\circ$
Wu et al. [25]
2023
$\circ$
$\circ$
Our survey
2024
✓: Paper covers the topic fully for Edge Learning; $\circ$ : Paper covers the topic partially for Edge Learning; •: Paper covers the topic with a focus on both Edge Learning and Inference; \: Paper covers the topic only from an Edge Inference point of view; ✗: Paper does not cover the topic;
IV Requirements for edge learning
The successful training of ML models at the edge requires meeting specific requirements that dictate the efficacy and performance of these models. These requirements can be viewed as metrics or evaluations, and they are essential for ensuring that the models perform optimally and efficiently within the resource-constrained environment inherent to edge devices. While there is no single metric to define the efficiency of training ML models in the edge [26], different ones can be used to make an estimation of how well the model will perform. We define some of these metrics/requirements in the figure 2.
Figure 2: Requirements of Edge Learning
  1. 1.
    Computational Efficiency: Computational Efficiency refers to the ability of an algorithm to achieve high performance with minimal computational cost. This is especially important in the context of EL, as edge devices, have limited computational resources [7], especially mobile edge devices which are even more resource-poor relative to static devices [27, 28]; and the training of ML models typically require high computational complexity [29].
  2. 2.
    Memory Footprint Efficiency: Similarly to computational complexity, edge devices often have low memory availability [30, 31], in contrast with the large memory footprint of ML models. This makes, memory efficiency an important requirement for EL.
  3. 3.
    Fast Training Time: Fast Training Time pertains to the rapid convergence of model parameters during the training phase. The importance of fast training time lies in its direct impact on the efficiency and responsiveness of edge devices. Edge devices, often characterized by limited computational capabilities, require ML models to be trained swiftly to minimize the processing burden and reduce energy consumption. Achieving Fast Training Time also ensures that models can be quickly adapted to evolving data patterns and ensures responsiveness and adaptivity where models can be efficiently updated to address dynamic environments and changing user requirements.
  4. 4.
    Optimized Bandwidth: Optimized Bandwidth refers to the optimization of the amount of data transferred in and out from edge devices, and the improvement of communication efficiency between edge devices. Bandwidth optimization is particularly important for distributed learning techniques at the edge and especially in bandwidth-limited systems, since these techniques require frequent sharing of the ML model across the network devices.
  5. 5.
    Low Energy Consumption: Energy consumption for edge devices is a crucial consideration, especially in mobile edge computing, considering the limited energy available at such devices. For this reason, ML models trained at the edge, need to be energy efficient, allowing for better computing performance and battery lifetime while ensuring the successful training of the models. Energy efficiency refers to the ability to use less energy to perform a specific task or function, it involves reducing energy waste and optimizing energy consumption.
  6. 6.
    Labelled Data Independency: The majority of the data produced at the edge are unlabelled data [32], for that reason, it might be beneficial to use in EL, ML techniques that are fully or partially trained using unlabelled data such as unsupervised learning (Section VI-A), self-supervised learning (Section VI-D) or semi-supervised learning (Section VI-C).
  7. 7.
    Task Specific Metrics and Performance: As Edge learning encompass different ML tasks and use-cases, specific metrics and benchmarks are used in general for the evaluation of the performances of the model and are vital for assessing its effectiveness in achieving its intended goals.
V Overview of Edge Learning techniques
Ultimately, the possibility of training an ML model on the edge depends on the resources requirements of the model, and the resources available at the device. With edge devices having increasingly more processing power, energy storage, and memory capacity available [33], it is currently possible to train a small ML model in an edge device without having to particularly optimize it for that regard or distribute the training across multiple devices. For examples KMeans in [34], Self-Organizing Map in [35] and SVM in [36]. However the training of more complex models that require heavier resources, such as neural networks is more challenging at the edge. Therefore, in this section, we will present an overview of techniques used to train more complex ML models in the edge, either through distributing the training to multiple devices, by training part of the model on the cloud and fine-tuning it on the local device or by optimizing or compressing the model to facilitate the training on the edge.
The Figure 3 shows a global view of Edge Learning techniques reviewed in this paper. The techniques are separated into four categories: (i) Distributed or collaborative techniques, such as federated or split learning; (ii) Techniques that rely on fine-tuning of a model trained on the cloud, such as Transfer or incremental learning; (iii) techniques that compress models to facilitate or support the training on the edge, such as Quantization and knowledge distillation; (iiii) And Finally the other Optimization Techniques that don’t fit needlessly into the previous categories.
{forest}
[…]
collaborative learning (V-A), text depth=3ex, text width=27ex, minimum height=1cm [Federated Learning (V-A1)] [Split Learning (V-A2)] [Swarm Learning (V-A3)] [Gossip Learning (V-A3)] ] [Adaptive and Fine-tuning
based Technique (V-B), text depth=3ex, text width=27ex, minimum height=1cm) [Transfer Learning (V-B1)] [Incremental Learning (V-B2)] [Meta Learning (V-B3)] ] [Model compression (V-C), text width=27ex, minimum height=1cm [Quantization (V-C2)] [Knowledge Distillation (V-C1)] [Model Pruning (V-C3)] ] [Other optimization
methods (V-D), text depth=3ex, text width=27ex, minimum height=1cm [Binary Neural Networks (V-D1)] [Spiking Neural Networks (V-D2)] [Forward-Forward Algorithm (V-D3)] ] ]
Figure 3: An overview of Edge Learning techniques
V-A Distributed and Collaborative Techniques
In this section, we will explore distributed techniques to train machine learning models at the edge. They work by leveraging the computational capabilities of multiple edge devices, and aggregating their results, instead of relying on a single resource constrained device.
V-A1 Federated Learning
Federated learning (FL), offers a transformative approach to decentralized model training. In the context of edge learning, where data is distributed across numerous edge devices, FL enables collaborative training without centralizing sensitive data [37]. This technique involves training a shared model across these devices by iteratively updating it based on local data, while preserving data privacy and minimizing communication overhead. FL is used extensively for training ML models at the edge [9], and has been adopted for multiple use-cases that ranges from detecting cyberattacks [38, 39] to spam detection [40, 41], or used in smart cities [42, 43] and autonomous vehicles [44]. Overall, privacy preservation is usually the main objective for FL [45, 46, 47].
In order to train an FL algorithm, an aggregation method is needed, Federated Learning aims to generate a global model by aggregating local models from multiple clients, this process combines the individual models to create a generalized model that represents the collective knowledge of all clients. The main two aggregation methods being, FedSGD (Federated Stochastic Gradient Descent) and FedAVG (Federated Averaging) [37]. However, other approaches have been proposed over the years such as EdgeFed [48] that reduce the number of calculations and time consumption of FedAvg by separating the process of updating the local model that is supposed to be completed independently by mobile devices. FedSel [49] an alternative to FedSGD that aims to solve its dimension dependency problem, by selecting Top-k dimensions according to their contributions in each iteration of federated SGD. Other approaches include MTFeeL [50], FedDynamic [51], FedNets [52], FedCom [53], FedGPO [54], and FedOVA [55].
Despite its growing popularity and multiple benefits, traditional FL models suffer from some limitations, for instance Non-IID (non-independent and identically distributed) data, often negatively impacts the performance of the global model [51]. FL is also susceptible to malicious and low-quality users [56], emerging new classes with completely unseen data distributions from any edge device, whose data cannot be accessed by the global server or other users [57] as well as single node failure [58, 59], channel bandwidth bottlenecks [60], and scaling issues for increasing network size [58]. To solve the low performance with Non-IID data challenge, Hybrid Federated Learning approaches (Hybrid FL) have been proposed [61], where very small amounts of data is shared from the participants. Other approaches that aim to solve this problem include FedNets [52], FedDynamic [51] and [62] that proposes a one-shot Neural Architecture Search technique. Pairwise Correlated Agreement (PCA) [56], on the other hand, is a method that aims to evaluate individual users’ contribution to avoid malicious and low-quality contributions from users. Sharma et. al [63] proposes a framework to study different noise patterns in user feedback, and explore noise-robust mitigation techniques for training FL models, and [57] propose a unified zero-shot framework to handle emerging classes in edge devices.
Over-The-Air Federated Learning (OTA-FL) [64] is another concept in FL that optimizes wireless channels for simultaneous edge device transmission of model updates, aggregating the models over the air in a ”one-time” manner, and minimizing resource consumption, this approach enhances communication efficiency through shared spectral resources and achieve efficient and low-latency global model aggregation [65]. Multiple approaches using OTA-FL have been developed in the past years. Mital and Gunduz [66] proposes an over-the-air computation (OAC) scheme for block-fading MACs that modulate the source to fit the available channel bandwidth in a wideband channel. [67] evaluates the system performance of FEEL with misaligned OAC. [68] propose a Hierarchical Over-The-Air Federated Learning (HOTAFL) method to mitigate the low performances of OTA-FL by the presence of mobile users located far away from the server. Other OTA-FL approaches include [69, 70, 71, 72].
Hierarchical Federated Learning (HFL) and Hierarchical Federated Edge Learning (HFEL) are an extension of FL that introduces a multi-level architecture [73]. These architectures allow for more efficient communication and computation trade-offs [74], faster model training, and reduced energy consumption by offloading tasks to edge servers for partial model aggregation to reduce network traffic [75]. [76] propose a H-FEEL training algorithms that address challenges in helper scheduling and communication resource allocation, and [75] propose a task offloading approach based on data and resource heterogeneity to improve training performance and reduce system cost. Other variations of FL/FEEL include Blind Federated Edge Learning [77], Modular Federated Learning [78], and Clustered Federated Learning (CFL) [79] that will be presented in more details in section VI-A.
Training language and multimedia models at the edge using FL is increasingly explored. While the training of Large Language Models (LLMs) using FL is still experimental, some approaches have been proposed such as FATE-LLM [80] and FwdLLM [81] that aim to fine-tune billion-sized language models across mobile devices using FL. On the other hand, there has been more work on the training of smaller language models such as BERT [82] with FL, such as FedBERT [83] that uses FL and SL approaches for pre-training BERT in a federated way, while FedSplitBERT [84] handles heterogeneous data and decreases the communication cost by splitting BERT encoder layers into a local part trained locally by the client and a global part trained by aggregating gradients of multiple clients. On the other hand, FedSPAM [40], fine tune a distilBERT model [85] using FL on mobile devices to detect spams in SMSes. In computer vision, FedVKD [86] was proposed as a federated knowledge distillation training algorithm to train small CNN models on edge devices and periodically transfers their knowledge to a large server-side vision transformer encoder via knowledge distillation, and [87] introduces an FL approach for visual classification with real-world data distribution. Finally, training audio models at the edge using FL is wildly explored for tasks such as speech recognition [88, 89, 90, 91, 92] or audio classification [57, 93].
Federated Learning is also used alongside other techniques presented in this survey, such as split learning [94], meta learning [95, 96], transfer learning [97], knowledge distillation [25, 98, 99], and Quantization [100, 99, 84], etc.
V-A2 Split Learning
Split Learning (SL) is an alternative collaborative learning method that enables clients to train ML models without sharing their data. In contrast to FL, where models are trained on local data from different devices then aggregated on a central server, SL divides the model into sections. Each section is trained on a different client. Instead of transferring raw data, only the weights of the last layer of each section are sent to the next client. This process ensures model improvement while maintaining better data and model privacy than FL due to the model architecture split between clients and the server, the split also makes SL a better option for resource-constrained environments, at the cost of slower processing than FL due to its relay-based training [94].
Split Learning at the edge has been increasingly studied in the last couple of years (see figure 5), such studies include, SplitEasy [101] a framework for training ML models on mobile devices using SL, [102] that proposes a data protection approach for SL without compromising the model accuracy, and [103] proposes a split edge learning based method by designing an online model splitting and resource provisioning game scheme that aims to minimize the total time cost of participating devices. Adaptive SL (AdaSplit), is a branch of SL that aims to solve its shortcomings compared to FL (slower processing, suboptimal performance and high bandwidth consumption), AdaSplit addresses these challenges by eliminating the transmission of gradients from the server to the client, resulting in a smaller payload and reduced communication cost, and allowing the client to update only sparse partitions of the server model, adapting to the variable resource budgets of different clients which decreases the computation cost and improves performance across heterogeneous clients [104, 105]. Other AdaSplit approaches for edge learning include ARES [106] and [107]. Finally, SL has been combined with FL multiple times in order to eliminate both techniques’ inherent drawbacks, notably in [94, 108, 109, 110, 111, 86, 112].
V-A3 Other Collaborative Learning methods
Several distributed learning techniques have been proposed as alternatives to FL and SL. Swarm learning, an innovative approach integrating artificial and biological intelligence, addresses challenges in distributed ML for the edge. This method efficiently utilizes signal processing and communication techniques to operate in real-time within large-scale edge IoT environments, offering advantages in overcoming communication bottlenecks, diverse data, non-convex optimization, and privacy concerns [113]. As another approach to Swarm learning, CB-DSL [114], a Communication-efficient and Byzantine-robust Distributed Swarm Learning technique, was introduced to deal with non-i.i.d. data issues and Byzantine Attacks. Another noteworthy distributed learning method is Gossip Learning, which, like other collaborative methods, doesn’t require transferring data outside edge devices. However, unlike FL and other methods, Gossip Learning operates without a central server for model aggregation and lacks reliance on central control [115]. Notable extensions to Gossip Learning, such as the one proposed by [116], enhance the algorithm by incorporating additional memory for storing local caches of model updates, making it more suitable for mobile devices.
V-B Adaptive and Fine-tuning based Techniques
In this section, we discuss techniques for efficiently adapting and fine-tuning pre-trained ML models at the edge without requiring complete retraining. The focus of these techniques is on preserving privacy and achieving personalized performance, while reducing computational overhead by keeping the heaviest part of the training in the cloud or edge servers, either using public datasets or ethically collected data. We explore three approaches: transfer learning, incremental learning, and meta learning. These methods enable edge devices to leverage previously acquired knowledge, adapt to local data distributions, and the continuous improvement of the models on the edge.
V-B1 Transfer Learning
Transfer Learning (TL) is an ML technique where knowledge gained from solving one problem is applied to a different, yet related, problem. Instead of building models from scratch, TL employs pre-trained models on large datasets to extract valuable insights, such as learned features or representations. These insights are then used to enhance the performance of a new task, especially when limited data is available for that task. By capitalizing on existing knowledge, TL accelerates model training, improves generalization, and proves exceptionally useful in domains where data scarcity poses a challenge [117].
In the context of edge learning, TL is a prominent technique used to fine-tune ML models based on local data in an edge device, and can serves as a fully on-device alternative to collaborative learning methods such as FL or SL that distribute the training across different devices [31], here are some state-of-the-art methods for transfer learning techniques for edge learning. TinyTL (Tiny-Transfer-Learning) [30] aims to achieve on-device learning memory efficiency, which is especially important for low memory edge devices. It achieves that by freezing the weights of the model and only learning a memory-efficient bias module, thus removing the need to store the intermediate activations. Similarly, RepNet [118] proposes an intermediate feature re-programming of a pre-trained model with a tiny Reprogramming Network to develop memory-efficient on-device transfer learning, and MobileTL [119] proposes a memory and computationally efficient on-device transfer learning method for models built with Inverted Residual Blocks. [120] propose an edge CNN framework for 5G industrial edge networks, with the CNN model trained in advance in an edge server, which is further fine-tuned based on the limited datasets uploaded from the devices with the aid of Transfer Learning. And [121] proposes a runtime convergence monitor to achieve massive computational savings in the practical on-device training workloads. Multiple approaches also focus on combining TL and FL, to create federated transfer learning algorithms [122], that aim to leverage FL for privacy preservation, and use transfer learning to train a well-performing local model despite users usually having not enough data for that by training the base model with a public dataset and passing it to the federated users to be fine-tuned for the target task [97, 123]. Finally, [124] proposes freeze and reconfigure, a TL method for on-device training of a BERT model.
V-B2 Incremental learning
Incremental learning also referred to as continual learning or life-long learning, involves continuously updating and expanding a model’s knowledge as new data becomes available. Unlike traditional batch learning, where models are trained from scratch on entire datasets, incremental learning dynamically incorporates new information without discarding previously acquired knowledge [125] [126], and can be used to reduce/overcome the well-known issue of catastrophic forgetting in deep neural networks [127, 128, 129].
There has been considerable attempts of implementing incremental learning in the context of edge learning. These include: learning with sharing [130] which aims to reduce the training complexity and memory requirements while achieving high accuracy during the incremental learning process and bypass the considerable memory requirements that can make incremental learning unsuited for edge devices; PILOTE [127] that trains an incremental learning model on edge devices for human activity recognition; [131] introduces an incremental algorithm based on transfer learning and k-nearest neighbor to support the on-device learning; RIANN [132] is an indexing and search system for graph-based approximate nearest neighbor algorithm for mobile devices; and RILOD [133] which aims to incrementally train an existing object detection model to detect new object classes without losing its capability to detect old classes, to avoid catastrophic forgetting. RILOD distils three types of knowledge from the old model to mimic the old model’s behaviour on object classification, bounding box regression and feature extraction, and it was implemented under both edge-cloud and edge-only setups [133]. There are a variety of promising approaches and directions for incremental learning on the edge from combining it with other techniques (such as Federated learning [134, 135], meta learning [136, 135] and compression methods [137, 138, 139]) to Sparse [140] or Distributed Continual Learning [138].
V-B3 Meta-learning
Meta learning focus on enhancing a model’s ability to learn new tasks quickly and effectively. Unlike traditional learning paradigms that optimize for a specific task, meta learning trains models to learn from a diverse set of tasks, thereby enabling them to generalize knowledge and adapt rapidly to novel tasks with minimal data [141]. By exposing models to various learning scenarios, meta learning equips them with transferable skills, such as recognizing patterns and adapting to new contexts.
In the context of edge learning, the application of meta learning introduces a transformative approach to address the challenges posed by limited data availability and resource constraints [142, 143]. [144] proposes adaptation-aware network pruning, a model pruning method designed to work with existing meta learning methods to achieve fast adaptation on edge devices, while [136] proposes a continual meta-Learning approach with Bayesian Graph Neural Networks that mathematically formulates meta-learning as continual learning of a sequence of tasks, and p-Meta was introduced in [142], and aims to achieve faster generalization to unseen tasks and enforces structure-wise partial parameter updates to support memory-efficient adaptation. Meta Learning can also be used in accordance to other techniques such as federated learning [135, 145], or [95] that integrates reinforcement learning models trained by multiple edge devices into a general model based on a meta-learning approach, in order to create FedMC, a generalized federated reinforcement learning framework based on a meta-learning approach.
V-C Model Compression based Techniques
Within this section, we embark on an exploration of model compression techniques aimed to streamline the training of ML models at the edge. As traditional deployment and inference solutions have embraced knowledge distillation, quantization, and model pruning to streamline model execution on resource-constrained devices, a notable shift is observed towards employing these techniques for reducing the complexity of ML models for the training in the edge, making these techniques helpful for the training phase as well.
V-C1 Knowledge distillation
Knowledge Distillation (KD) in deep learning is a process whereby a small or student neural network is trained to emulate the knowledge and predictive capabilities of a larger or teacher network. This technique serves as a means to transfer the expertise and generalization capabilities of a complex model to a simpler one, thereby enhancing inference efficiency and reducing computational demands [146]. The underlying principle involves the student network learning not only from ground truth labels but also from the soft, probabilistic outputs of the teacher network, thereby capturing finer details and nuances in the data [146]. In the context of edge learning, knowledge distillation is usually used to reduce the size and complexity of a large neural network, to simplify its training in limited resources devices. Therefore, knowledge distillation is well suited to be used in collaboration with other techniques such as federated learning [25, 98], split learning [147] or incremental learning [137]. However, distillation is also used independently of other techniques [148, 149].
In the context of using knowledge distillation with FL on edge devices, recent trends indicate great potential in combining the two techniques [25], and considerable approaches have been developed in that regard, from an Attack-Resistant Federated Edge Learning method with dataset distillation [150] to speech recognition tasks [92] or keyword spotting [151]. Mix2FLD [152] is another method that combine KD and FL, while [99] use both KD and Quantization to train Federated Learning models on Edge Devices. Other hybrid methods involving FL and KD have been proposed [153, 154], some of them are explored in details in Wu and Al’s Survey of Knowledge Distillation in Federated Edge Learning [25]. Other Distributed learning methods can benefit from KD as well, for instance [147] introduces a spatio-temporal distillation method for split learning for a tiny server in order to alleviate the frequent communication costs that happen when communicating from the server to edge devices, and [155] introduces a distributed distillation algorithm where devices communicate and learn from soft-decision outputs, which are inherently architecture-agnostic and scale only with the number of classes in order to alleviate the communication costs from transmitting model weights in the network and improve the inclusion of devices with different model architectures. Finally, KD has been used as a standalone technique in an edge learning context, for recommendation systems [156, 157, 148], Edge Cardiac Disease Detection [158] and on-device Deep Reinforcement Learning [159]. KD was, additionally, used with multiple variants, including dataset distillation techniques [160, 161] and knowledge transfer [162, 149].
V-C2 Quantization
Quantization in deep learning refers to the process of reducing the precision of numerical values representing model parameters or activations, typically from floating-point to fixed-point or integer representations, in order to balance the act of maintaining an acceptable level of model accuracy while significantly reducing the memory and computational requirements [163]. This computational optimization technique is pivotal in mitigating the resource-intensive demands of deep neural networks, rendering them more amenable for resource-constrained hardware platforms, such as edge devices and embedded systems[163]. There are two types of quantization: Quantization-Aware Training (QAT) and Post-Training Quantization (PTQ). In QAT, the quantized model is fine-tuned using training data in order to adjust parameters and recover accuracy degradation or perturbation introduced by the quantization, while in the less expensive PTQ the pretrained model is calibrated using calibration data before the quantization which is based on the calibration result [163].
Although quantization techniques are usually used to optimize machine learning models before deployment for edge inference [164, 121], it can be used in edge learning in order to simplify the fine-tuning of large models in the edge [31, 165]. Similarly to KD, quantization is often alongside other techniques such as federated learning [166], transfer Learning [31], incremental learning [167] or with other types of techniques [168]. Quantization is used with FL particularly extensively, from one-bit quantization [169, 170, 171] to hierarchical FL [100] and other FL based approaches such as [172, 173, 174, 175, 176, 99]. Other Quantization based methods for training ML models in the edge include, Quantization-Aware Scaling (QAS), a method proposed in [31], to automatically scale the gradient of tensors with different bit-precisions without requiring any fine-tuning, the method was used alongside the Tiny Training Engine and Sparse Update introduced in the same paper [31]. In [177], investigations were made on on-device few shot learning for audio classification, and it was found that quantization helps further in reducing the resource requirements for the training. In the optimizer. Holmes [178], quantization is used in an ingenious way to improve the accuracy by effectively combining different types of quantization (a limitation on the number of quantization bits, fixed-point number, and logarithmic quantization) techniques.
V-C3 Model Pruning
Model pruning is a technique used to reduce the size of ML models by removing certain parts of the model, such as model parameters, nodes in a decision tree [179] or weight matrices in transformer-based models [180]. Similarly to quantization, model pruning is commonly used in the edge to improve the efficiency of inference and reduce computational resources required for the deployment of ML models [181]. However, much like quantization, model pruning has also shown great potential in edge learning, by reducing the size of the ML model before fine-tuning it on edge devices, particularly used in conjunction with other techniques such as Federated learning [174, 182, 183, 184], Incremental learning [131, 139] or meta-learning [144].
Similar to knowledge distillation and quantization, federated learning emerges as the most prominent technique when combined with model pruning for EL. Noteworthy is PruneFL [183], an approach aimed at minimizing communication and computation overhead while reducing training time through adaptive model size adjustment during FL. PruneFL employs model pruning, starting with an initial pruning stage at a selected client, followed by subsequent pruning iterations during FL. Additionally, [184] introduce model pruning for wireless Federated Learning to scale down neural networks, while [182] employ an adaptive dynamic pruning approach to prevent overfitting by slimming the model through the dropout of unimportant parameters. Other approaches that use model pruning on the edge include [185] that uses model pruning in the context of On-Device Personalization for an activity recognition system, and Deeprec [186] which leverages model pruning and embedding sparsity techniques to reduce computation and network overhead. OmniDRL [187], is a Deep Reinforcement Learning based approach on edge devices, incorporates weight pruning in each learning iteration to achieve a high weight compression ratio. Finally, [188] explores the reduction in memory footprint for further pruning during the training phase of BitTrain, a bitmap memory efficient compression technique for training on edge devices.
V-D Optimization and Acceleration based Techniques
In this section, we explore some of the multiple other techniques that do not necessarily fall into a specific category and are used to optimize or provide more optimized alternatives to machine learning models’ computation efficiency, memory footprint, energy consumption, etc, which makes them more suited for edge learning.
V-D1 Binary neural networks
Binary Neural Networks (BNNs) are Deep Neural Networkss (DNNs) that use binary values (-1 or 1) instead of floating-point numbers for weights and activations. BNNs are attractive for resource-constrained devices due to their ability to compress DNNs [189]. Similar to other techniques such as quantization and model pruning, BNNs are considered to be good candidates for edge inference due to their extreme compute and memory savings over higher-precision alternatives [190]. However, multiple approaches were proposed to leverage BNNs compute and memory efficiency for edge learning, such as [191] that proposes a hybrid quantization of a Continual Learning model with a BNN as its backbone, or [192] that propose a BNN model based on an MRAM array with Ternary Gradients for both training and inference on the edge, other BNN based approaches for edge learning include [193, 194, 190].
V-D2 Spiking neural networks
Spiking neural networks (SNNs) are another type of DNNs that is promising for the edge, SNNs communicate between neurons using events called spikes [195] and are known for their asynchronous and sparse computations, which result in decreased energy consumption [196, 197], and this make them well suited for edge computing especially for energy limited devices [198]. Training ML models at the edge using SNNs had gathered some traction, [199] propose FL-SNN, a cooperative training through FL for networked on-device SNNs, [200] present a memristor spiking neuron and synaptic trace circuits for efficient on device learning, and [143] integrate a meta-learning approaches with SNNs that enables lifelong learning on a stream of tasks, and relies on local backpropagation-free nested updates. SNNs can also be trained at the edge using event-driven, power and memory-efficient local learning rules, such as spike-timing-dependent plasticity (STDP), such as [201] that use STDP and Synaptic Normalization for unsupervised on-device learning. There are other approaches that leverage SNNs for EL, including [202, 203, 204].
V-D3 Forward-Forward Algorithm
The backpropagation algorithm is essential for training neural networks, but recent studies have proposed alternatives to the algorithm when the available resources are limited, such as the Forward-Forward (FF) [205] algorithm that aims to replace the forward and backward passes of backpropagation by two forward passes that operate in the same way as each other on different data and with opposite objectives. A positive pass operates on real data and adjusts the weights to increase the goodness in every hidden layer, while a negative pass operates on ”negative data” and adjusts the weights to decrease the goodness in every hidden layer [205]. Some other studies have tried to adapt the Forward-Forward algorithm to edge devices, such as µ-FF [206] that propose an on-device variation of the original FF on microcontroller units which tackles the training process with a multivariate Ridge regression approach and allows finding closed-form solution by using the Mean Squared Error (MSE) as loss function. And finally, [207] investigates the improvements in terms of complexity and memory usage brought by PEPITA [208] and the Forward-Forward Algorithm, the results show that the Forward-Forward algorithm reduce memory consumption by 40% on average, however it involves additional computation at inference that, can be costly on microcontrollers.
V-D4 Other techniques
In this section, we will explore some other techniques used to optimize ML models for training in the edge. One such techniques is Data Booleanization used in [209] which proposes a novel approach towards low-energy booleanization. MiniLearn [210] on the other hand, enables re-training of DNNs on resource-constrained IoT devices and allows IoT devices to re-train and optimize pre-trained, quantized neural networks using IoT data collected during the deployment of an IoT device. While [211], propose the use of Echo State Networks for anomaly detection on-the-edge in aerospace applications. Tiny Training Engine (TTE) [31] is a lightweight training system, introduced alongside Sparse Update, a technique that skip the gradient computation of less important layers and sub-tensors, and a Quantization-Aware Scaling to stabilize 8-bit quantized training, TTE enables on-device training of convolutional neural networks under 256KB of SRAM and 1MB Flash without auxiliary memory [31].
[212] introduces a novel reduced precision optimization technique for On-Device Learning primitives on MCU-class devices with a specialized shape transform operators and matrix multiplication kernels, accelerated with parallelization and loop unrolling for the backpropagation algorithm. POET [213], allow for the training of large neural networks on memory scarce and battery-operated edge devices, with Integrated Rematerialization and Paging POET reduce the memory consumption of backpropagation and could fine-tune both ResNet-18 and BERT within edge devices memory constraints. Finally [168] proposes a novel rank-adaptive tensor-based tensorized neural network mode for on-device trainging with ultra-low memory usage.
V-E Comparison of techniques used in edge learning
In this section we will compare the different families of techniques used to train ML models in the edge that we explored in the previous part of this paper (sections V-A, V-B, V-C, V-D), we will compare the different families based on two factors : 1. The number of academic contributions of each different technique family and their evolution over the years; 2. The techniques’ potential of answering the different needs and requirements particular with edge learning.
V-E1 Comparison of the usage of the different techniques
We will first start by making a comparison of the different EL techniques’ over the years, by analysing their academic contributions. The figure 4 present the number of papers for each technique, while the figure 5 references the number of papers (in a logarithmic scale) per technique per year. In order to get the necessary data, We used the advanced search feature of Scopus  [^1] Scopus : https://www.scopus.com/ where we searched the terms ”edge learning”, ”training/learning on the edge/mobile devices”, ”on-device training/learning” and ”on-device adaptation” as well as the relevant keywords for each technique (”Federated Learning”, ”Split Learning”, etc), in the Title, Keywords and Abstract, we made sure to exclude the surveys, conference reviews, books and notes as we are only interested in approaches to train ML models in the edge using the aforementioned techniques. For each family of techniques, we manually reviewed and removed papers that either didn’t provide a contribution, or weren’t about training ML models on edge devices with said technique, despite checking the keywords. Finally, we added a few papers we found manually and either were not available in Scopus or didn’t check the keywords but were relevant in the contest of the analysis. The final number of papers associated with the analysis was 511 papers, and some of these papers were briefly covered in V, note that multiple techniques can be used in a single paper, therefore the total count of techniques in the figures 4 and 5 will exceed 511.
Figure 4: Overall trend of techniques used to train ML models in the edge
Figure 5: Trend of technique used to train ML models in the edge over tthe years
The analysis of Figure 4 reveals that Federated Learning (FL) is the dominant approach for training ML models in edge environments, given the resource constraints of edge devices. This dominance is expected, as distributed learning methods that capitalize on the collective computing power of multiple devices are deemed more practical and efficient in edge settings. Moreover, we anticipate that this trend will persist and expand to include Split Learning (SL), another promising distributed learning technique. Other methods, including Incremental Learning (IL), Transfer Learning (TL), Model Compression Techniques (e.g., Quantization, Knowledge Distillation (KD)), although consistently employed, lag behind FL in terms of popularity. Furthermore, techniques such as Swarm or Gossip Learning have received scant attention. A closer examination of Figure 5 unveils a remarkable surge in the number of publications focused on edge learning over the past six years. Notably, there has been a steady rise in the adoption of FL and Split Learning from 2018 to 2023, aligning with our forecast of a burgeoning trend favoring these two techniques. In contrast, the use of techniques like IL and TL has been more steady during the same period, and meta-learning despite being employed consistently from 2019 to 2022 received little attention in 2023. Finally, for the model compression techniques while quantization and KD experienced a small rise in popularity over the years, model pruning has exhibited greater fluctuation during that period.
As previously discussed, various approaches have been proposed that integrate multiple techniques to mitigate the limitations of individual methods and capitalize on their respective strengths. Examples of such approaches include [94, 135, 136, 122]. To provide a clearer illustration of the relationships between these various techniques, a heatmap depicting the intersection of their usage is presented in Figure 6. This visual representation allows for a more comprehensive understanding of the synergies and overlap between different approaches. We can note that FL has been combined the most with other techniques, which is expected considering the overwhelming number of FL contributions to the edge (see Figure 4). Additionally, Model Compression techniques such as KD, model pruning and quantization are also used together with other techniques as discussed in Sections (V-C1 V-C2 V-C3). Finally, we can note other collaborations between the different families, and we expect this trend to continue in the future.
Figure 6: Trend of different techniques used hand in hand for training ML models at the edge. Color intensity represents the number of papers
V-E2 Comparison based on the requirements and needs of EL
As discussed in Section IV, several requirements must be met for edge learning, which can function as imprecise measures for assessing the viability of the families of techniques covered in earlier sections. After reevaluating the prerequisites, we eliminated ”Labelled Data Independence” from the comparison, because it is more relevant to the type of machine learning (ML) employed and the availability of an autolabeling process than to the approaches under consideration. Additionally, we do not consider ”Task-specific metrics and performance” since that covers multiple metrics used in specific tasks; nevertheless, we introduce a ”High Performance” measure to estimate roughly if the strategies positively or negatively affect the performances. For instance, model compression techniques often lower model performance, whereas incremental and meta-learning typically boost it. Consequently, we employ six distinct measures: ”Computational Efficiency,” ”Memory Footprint,” ”Low Energy Consumption,” ”Quick Training Time,” ”Optimized Bandwidth,” and ”High Performance.” Table II compares the various families of techniques against these requirements.
The outcomes, as depicted in Table II, should be interpreted as informal and approximate assessments aimed at providing a broad understanding of the general strengths and weaknesses of the compared techniques within the context of edge learning (EL). A checkmark (”✓”) implies that, in general, the technique offers assistance or advantages when applied to EL with respect to the specified requirement. Conversely, a cross mark (”✗”) indicates that, in general, the requirement represents a weakness of the technique in the EL context. It is important to note that techniques like Federated Learning (FL), Quantization, and others have various variants and specific approaches that influence how these methods align with the requirements. For instance, while Quantization techniques may result in a minor decrease in performance in most scenarios [214], leading to an ”✗” check in the ”High Performance” column. Certain specific approaches to Quantization may exceptionally yield no performance loss or even improvements, as demonstrated in studies such as [178]. Therefore, it is important to note that the assessment provided by these symbols should be considered as rough estimates, as the effectiveness of a technique, and how it fares against a specific requirement, can vary depending on diverse factors such as variant versions, implementation details, use cases, tasks, and hardware platforms. Accordingly, the table II offers a general overview rather than a definitive judgment on the suitability of each technique for every situation.
TABLE II: Comparison between The different techniques that enable edge learning
Technique Computation Efficiency Memory footprint Low energy consumption Fast Training time Optimized Bandwidth High Performance
Federated Learning
Split Learning
Swarm Learning
Gossip Learning
Transfer Learning
Incremental Learning
Meta-Learning
Knowledge Distillation
Quantization
Model Pruning
BNNs
SNNs
Forward-Forward Algorithm
✓: Have a positive effect on the requirement;
✗: Have a negative effect on the requirement;
•: Have a neutral or uncertain effect based on specific conditions on the requirement;
\ : There is not enough information and literature to estimate the effect on the requirement
VI Edge Learning for different types of Machine Learning
In this section, we will explore the usage of different types of ML in the edge. We will focus on unsupervised learning, reinforcement learning, semi-supervised and self-supervised learning, as they present some particularities when adapted to the edge, however, we will ignore supervised learning [215] as it’s usually considered the default when it comes to model training and most approaches at the edge use it without requiring any adaptation or particular implementation.
VI-A Unsupervised Learning
Considering the vast amount of unlabeled data produced in edge and end devices [32], it is very promising to use unlabeled data to train ML models on the edge. However, unsupervised learning comes with multiple challenges and restrictions for edge learning, especially when it comes to collaborative learning techniques such as FL or SL, which represent the vast majority of techniques used in the edge (see Figure 4). Unsupervised learning datasets may have a non-identically and independently distributed (non-IID) nature. Each node in a collaborative setting might have a different subset of the data, and the data distribution might vary across nodes. This non-IID property can make it difficult to effectively combine information from different nodes. Additionally, in the case of clustering algorithms, clusters may have varying sizes across nodes in a collaborative setting and clustering algorithms may need to adapt to changes in data distribution and cluster structures over time. Finally, since there are no available labels, and their assignment may differ between nodes (for example in a clustering algorithm). Ensuring consistency across distributed nodes is difficult, but crucial for aggregating meaningful global labels (clusters).
The figure 7 represent the main different types of unsupervised learning approaches used for training ML models in the edge, (a) using unsupervised learning algorithms directly on-device with non-collaborative methods [216, 35]; (b) using unsupervised learning methods to assist in the training of collaborative learning approaches, such as clustered federated learning [79, 217, 218] ; and (c) training an unsupervised learning model on the edge in a collaborative [219].
(a) Unsupervised learning On a single device
(b) Unsupervised learning to assist collaborative learning
(c) collaborative unsupervised learning
Figure 7: Different types of Unsupervised Learning for Edge Learning
Unsupervised learning on a single edge device
Having an unsupervised learning trained on a single edge device is possible if the edge device has enough computation power and/or the learning algorithm is lightweight and can be trained with low resources, the training with such models is no different from the training on the cloud or other devices, the only difference being the constraint of low resources and data available [35]. Examples of unsupervised learning algorithms trained with this approach include [34] that investigate the application of K-Means on mainstream controllers, and [220] that presents the first dedicated CycleGAN accelerator for energy-constrained mobile applications, achieving a higher throughput-to-area ratio and higher energy efficiency than a GPU. In [216], an unsupervised segmentation was proposed that can be executed on edge devices without the need of annotated data. While [221], proposes TMNet, an approach to solve unsupervised video object segmentation problem at the edge. Finally, [222] propose an FPGA based architecture for a Self-Organization Neural Network capable of performing unsupervised learning on input features from a CNN by dynamically growing neurons and connections in order to perform class-incremental lifelong learning for object classification in the edge.
Unsupervised learning to assist collaborative learning approaches
Collaborative learning approaches like FL are promising solution for training ML models in the edge, However FL the most popular technique on the edge, face multiple challenges such as non-independent and identically collaborative (non-IID) data and uneven computing power [110] and yields suboptimal results when the local clients’ data distributions diverge [79]. To solve these issues, the usage of clustering alongside FL have been proposed multiple times to cluster edge servers with similar environmental data distributions [223]. Among these approaches, Clustered Federated Learning (CFL), introduced by [79] is widely adopted, and it works by exploiting geometric properties of the FL loss surface and grouping the client population into clusters with jointly trainable data distributions. Approaches that build upon CFL have also been proposed, and they include [217, 110, 218, 224]. Other approaches that differ from CFL [79] include [68] that propose a hierarchical over-the-air FL approach which utilizes intermediary servers to form clusters near mobile users; and HPFL-CN [223], a communication-efficient hierarchical personalized federated edge learning framework, that achieves this by utilizing complex network feature clustering to group together edge servers that exhibit similar environmental data distributions. Subsequently, personalized models are trained for each cluster using a hierarchical architecture, resulting in enhanced efficiency. More specifically, HPFL-CN incorporates Privacy-preserving Feature Clustering (PFC) to derive privacy-preserving low-dimensional feature representations for each edge server. This is achieved by mapping the environmental data onto various complex network domains, thereby accurately clustering edge servers with similar characteristics. And Finally, [57] uses unsupervised learning methods on the edge to distinguish between classes across different users, when new classes with completely unseen data distributions emerge on devices in a federated learning setting for audio classification.
Collaborative Unsupervised learning at the edge
as explained in previous sections, collaborative unsupervised learning methods are hard to train using unsupervised learning. Nevertheless, several methodologies have surfaced, one of which is FedUL (federation of unsupervised learning) [219]. In this approach, the unlabeled data undergo a transformation process to become surrogate labelled data for each client. Following this, a modified model is trained through supervised FL. Eventually, the desired model is obtained by recovering it from the modified model. An alternative methodology, known as FedUReID, has been proposed by Zhuang et al. [225] as a federated unsupervised person ReID system. The purpose of this system is to train person ReID models without the use of any labels, all the while ensuring the preservation of privacy.
VI-B Reinforcement learning
Reinforcement learning (RL) has been successfully applied in the past on different problems in areas such as robotics, recommendation systems, video games and automatic vehicles [226, 227], making RL a promising and interesting direction for edge learning. However, The training of RL models in resources constrained environments is often limited by high compute and memory requirements from gradient computations [228], making the application of RL in the context of edge learning challenging. Despite these challenges, multiple RL approaches have been proposed for training ML models in the edge. Among them, Federated Reinforcement Learning (FRL) [229, 230] is a promising approach that allow multiple RL agents to learn optimal control policies for a series of devices with slightly different dynamics [231]. FRL is employed to achieve diverse objectives including personalization [232, 233], IoT traffic management [234, 235], Autonomous Systems [231] and resource allocation for unmanned aerial vehicle (UAV) [236]. FedMC [95] integrates RL models trained by multiple edge devices into a general model based on a meta-learning approach. FedGPO is an RL-based aggregation technique for FL introduced in [54], and aims to optimizes the energy-efficiency of FL while guaranteeing model convergence. On the other hand, [237] introduces DDQN-Trust, a trust-based double deep Q-learning-based selection algorithm for FL that takes into account the trust scores and energy levels of the IoT devices to make appropriate scheduling decisions and integrate it with the main FL aggregation techniques (FedAvg, FedProx, FedShare and FedSGD). Other Federated Reinforcement Learning methods include [75, 238, 239, 240, 241, 242]
Other Reinforcement learning approaches that don’t rely on FL or other collaborative learning paradigms have been proposed, such as [243] that uses online sequential learning to achieve full on-device reinforcement learning on an FPGA platform. In [244], a method using a combination of supervised learning and reinforcement learning was proposed for adaptive video streaming on edge servers or on-device, and [245] introduces an on-device RL-based adaptive video transmission algorithm to predict heterogeneous network bandwidth. EdgeRL, a lightweight C/C++ RL framework, was introduced in [246] and aims to provide a platform to train RL on edge devices. Finally, RL has also been employed in edge learning via shielding techniques [247], as proposed in [248] with a multi-agent system that enables each edge node to schedule its own jobs using SROLE, a shielded reinforcement learning technique used to check for action collisions that may occur because of the absence of coordination between the nodes, and provides alternative actions to avoid them.
VI-C Semi-supervised learning
As mentioned in the section VI-A using unlabeled data to train on the edge is very promising considering the vast amount of unlabeled data in edge and end devices [32]. Semi-supervised learning, constituting a machine learning paradigm, seeks to leverage both labeled and unlabeled data for specific learning tasks [249], presenting a means to harness the vast amount of unlabeled data at the edge, circumventing the challenges associated with supervised and unsupervised learning. Semi-supervised learning also provides other advantage on the edge, as training datasets are often not completely defined before training and might need to be completed using real-time data [250]. Naturally, several methodologies have surface to adapt semi-supervised learning on the edge, using federated learning based approaches [251, 252], or other learning techniques [253, 254, 255, 211, 256, 250, 250, 257].
VI-D Self-supervised learning
Self-supervised learning (SSL) is a machine learning approach that allows models to learn from vast amounts of unlabeled data. SSL defines a pretext task based on unlabeled inputs, such as predicting the context surrounding a word in natural language or predicting masked patches of an image in computer vision, and learn generic representations that are useful across multiple tasks, and can be especially beneficial in domains where labelled data is scarce, or the specific task can not be known a priori [258].
For this reason, Self-supervised learning is considered as a promising approach to leverage the abundance of unlabeled data at the edge for various AI applications. By learning from unlabeled data without explicit labels, self-supervised learning models can learn useful representations and skills that can be fine-tuned for specific tasks, such as recommendation systems [156, 148], speech and audio-related applications [259, 260], and others [261, 262, 263, 264]. Moreover, self-supervised learning can be particularly effective in scenarios where obtaining high-quality labels is challenging or expensive and in situation with data and concept drifting [261]. Several recent studies have proposed innovative self-supervised learning methods tailored for edge devices, such as contrastive learning [261, 263, 264].
VII Edge Learning Use Cases and applications
As explained in previous sections, Edge Learning (EL) offer multiple advantages that range from low latency, bandwidth efficiency to privacy preservation and improved reliability and robustness, it also allows more customization and personalization by adapting to user preferences and behavior without relying on centralized computing or needing to collect and store private user data in cloud servers. In this section, we explore some uses cases and applications for EL that have been researched and developed in the past years.
VII-A Healthcare and Remote Monitoring
The use of ML in healthcare has been on constant improvement over the last years [265], despite this, cloud-based ML for healthcare continue to suffer from their limitations in meeting stringent security [266], privacy concerns [267], and low latency requirements [268]. As EL makes it possible to alleviate some of these challenges, it has gained considerable traction in the field of health care [45, 269, 270, 271]. Privacy concerns, being amongst the biggest challenges EL aims to solve for machine learning in healthcare [272, 45, 238]. Researches that aim to use EL for healthcare span in most of the field, from atrial fibrillation recognition [273], preterm labour risk prediction [274], cardiac disease detection [158], breast ultrasound image classification [255] to dermatological disease [264] and COVID-19 diagnosis, either using Clustered Federated Learning (CFL) [269] or blockchain technology and federated transfer learning on wearable devices [275].
VII-B Smart Technologies
Edge learning has emerged as a pivotal technological advancement for smart technologies such as smart cities [276], smart agriculture [277], smart homes [278], etc. In this section, we will explore some of its applications in these settings.
Smart cities are urban ecosystems designed using IoT technologies to solve urban life problems and improve residents quality of life [279]. In this context, edge learning has been proposed to solve different challenges, mainly for its ability to leverage ML capabilities while preserving network bandwidth and reducing the charge on cloud servers. In [280], a cloud-aided edge learning based on knowledge fusion for smart lighting system has been proposed. Another application of ML in the edge is in smart grid systems where ML is needed to improve demand forecasting and automated demand response, as well as to analyse data related to energy use and obtain energy consumption patterns [281], detect anomalies [282], improve communications [283] and security [284] in the system. Other applications in smart cities include the detection of abnormal and dangerous activities [285, 286], pedestrian detection [287], water consumption forcasting [42, 288] and reducing congestions in intelligent traffic systems [289, 290]. Other more general EL approaches for mart cities include [291, 292, 293, 276, 252].
Smart farming is another domain where ML is increasingly used to enhance the production quality, crop selection, and mineral deficiency detection, as well as to increase farmers’ earnings [294]. In [277] a TinyML based framework using deep neural networks and LSTM models for unmanned aerial vehicle (UAV) assisted smart farming was proposed, which measure soil moisture and ambient environmental conditions. Smart farming remains a promising domain for edge learning, although further research is needed for effectively harnessing its potential. Finally, using edge learning in smart homes can also be promising, however, at the time of writing this article, only a few papers explore this area, including [278].
VII-C Autonomous vehicles
Autonomous vehicles are vehicles that can operate without human intervention, they utilize sensor technologies, AI, and networking to navigate and make decisions [295]. Autonomous vehicles include self-driving cars trucks, buses, drones, unmanned aerial vehicles (UAVs), and even small robots. As demonstrated in [296] offloading deep learning tasks to edge devices or servers can improve the inference accuracy while meeting the latency constraint, which makes edge learning prefectly suitable for this use-cases, and as expected there has been extensive research done in this area.
Unmanned Aerial Vehicles (UAVs) are by far the most prominent use of EL for autonomous vehicles. In [297], a Synchronous Federated Learning (SFL) structure for multi-UAVs was proposed. That aims to resolve device privacy concerns that comes from sending raw data to UAV servers, as well as UAVs’ limited processing or communication resources. On the other hand, [298] propose a model-aided federated MARL algorithm to coordinate multiple UAVs on data harvesting missions with limited knowledge about the environment, significantly reducing the real-world training data demand. As mentioned previously in section VII-B, [277] aims to assist farming opperations using UAVs that measure soil moisture and ambient environmental conditions and [299] proposes a model to derive computation specifications for learning-based visual odometry (VO) from physical characteristics of UAVs. Other EL applications for UAVs include [236, 300, 301, 302, 303, 299, 304, 305, 306, 307]
Edge learning based approaches for other autonomous vehicles are also constantly explored and involve multiple applications, they include :
  • Trajectory predictions such as [308] that proposes a solution for trajectory prediction in the edge for both human-driven and autonomous vehicles by leveraging the capabilities of the 5G Multi-Access Edge Computing (5G-MEC) platform to collect and process measurements from vehicles and road infrastructure in edge servers and use a Long Short Term Memory (LSTM) model to predict the vehicle trajectory with high accuracy.
  • Energy efficiency for autonomous vehicles, where [309] proposes a rate-splitting multiple access (RSMA)-based Internet of Vehicles (IoV) system for energy-efficient Federated Edge Learning (FEEL) in autonomous driving, using Non-Orthogonal Unicasting and Multicasting (NOUM) transmission.
VII-D Recommendation systems and personalization
Recommender systems are intelligent applications that assist users in making decisions by providing advice on products or services they might be interested in [310]. However, recommender systems that utilize user data can pose threats to user privacy, such as the inadvertent leakage of data to untrusted parties or other users [311]. Furthermore, privacy-enhancing techniques may lead to decreased accuracy in the recommendations [312]. Edge learning, and especially collaborative learning approaches such as FL, have a big potential in solving these problems by allowing recommender models to be partially or completely trained on the edge, keeping user interactions on the device and using them to further personalize the system [313].
Different approaches using FL have been used for recommendation systems. Amongst them, FedFast [314] propose to accelerate distributed learning for deep federated recommendation models which achieve high accuracy early in the training process. In [315], a Graph Neural Networks (GNNs) used alongside federated learning for social recommendation tasks, a method that aims to alleviate the cold start problem by inducing information of social links between users [315]. On the other hand, [316] was proposed as a federated sequential recommender system for the edge. A method that, unlike traditional recommendations, provides personalized suggestions by sequentially analyzing users’ historical interactions [317]. To achieve this, [316] uses a knowledge-aware transformer and proposed to incorporate knowledge graph information into sequential recommendation tasks, while applying federated learning to preserve users’ privacy, and use replaced token detection and two-stream self-attention strategies to enhance the transformer-based model. Finally, FedCT [123] aims to harness cross-domain recommendation in the edge. While cross-domain recommendation [318] is a promising area for utilizing data from multiple domains, the conventional approach of sharing data between services in a cloud setting often proves impractical or impossible due to privacy and security concerns. This limitation emphasize the appeal of EL as an interesting direction for cross-domain recommendation. By enabling the training of recommender systems on multi-domain data residing on edge devices, while still respecting users’ privacy.
Despite the improvement of federated recommender systems for users’ privacy preservation [319, 320], distributed learning approaches for recommendation still face privacy challenges. Specifically, although users’ item ratings remain on-device, they can be inferred from the final model, thereby posing a risk of data leakage when the model is shared with multiple users [321, 322]. To address this concern, noise is often introduced to the ratings in the form of random user-item interactions. However, this approach usually results in lower performances [322]. In recent years, several solutions have emerged to mitigate this issue. For instance, FedMMF a federated masked matrix factorization, introduced in [321], aims to protect data privacy in federated recommender systems by using personalized mask generated only from local data. Another approach that aims to achieve that is FedRec++ [322], by allocating certain clients as denoising clients to eliminate noise in a privacy-aware manner, thereby counteracting the random sampling of items during the training phase. Other FL based edge recommendation systems includes [320, 323, 319, 324, 325]
Although the most popular approach, FL isn’t the only method to train recommender systems in the edge. In [186], an on-device deep learning sequential recommendation method aimed at mobile devices was proposed, by fine-tuning a pretrained model that was trained using data collected before GDPR  [^2] GDPR : The General Data Protection Regulation is a regulation on data privacy in the European Union and the European Economic Area, for further personalization. And [148] focus on on-device next-item recommendation, and uses compact models and a self-supervised knowledge distillation framework to compensate for the capacity loss caused by compression. Finally, [326] proposes a split-federated learning called SpFedRec where a split learning approach was proposed to migrate the item model from participant’s edge devices to the cloud side and compress item data while transmitting and apply a Squeeze-and-Excitation network mechanism on the backbone model to optimize the perception of dominant features.
Personalized crowdsourced livecast are another part of personalization methods that might benefit from being offloded to the edge. In [327] the rapid development of crowdsourced livecast and the challenges in providing personalized Quality of Experience (QoE) to viewers is discussed, and it introduces an intelligent edge-learning-based framework called ELCast, which integrates Convolutional Neural Networks (CNN) and deep reinforcement learning (DRL) in edge computing architectures for personalized crowdcast recommendation. In the area related to video game, personalisation involves constructing a system capable of adapting video game rules and content to better suit some aspect of the player preferences, personality, experience and performances [328]. Although not yet explored in edge and on-device learning, [329] propose a Deep Q network model to personalize games based on user-interaction on the edge.
VII-E Human Activity Recognition
Human activity recognition (HAR) refers to the automation of the identification and categorization of the various activities performed by humans and their interactions with the environment [330]. Since personalization for HAR has been shown to improved the results and performances of these systems [331], training the model on the edge using Incremental learning or meta learning approaches can help achieve that while increasing the privacy and reduce the bandwidth consumption. PILOTE [127] proposes an incremental learning based approach for HAR that is intended to be trained on edge devices with extremely limited resources and provides reliable performance in handling catastrophic forgetting. While [332] proposes an approach that proposes a personalizable lightweight CNN model for HAR. As well as a training algorithm to find personalization-friendly parameters, for boosting the accuracy after the personalization that is personalized and boost the HAR accuracy when dealing with a wide range of target users. ClusterFL [333], proposes a a clustering-based fl approach for HAR training on hte edge. Finally, [334] proposes an On-Device deep learning approach for STM32 microcontrollers by fine-tuning a 1-D CNN model for further HAR personalization.
VII-F Others
There are multiple other applications and use cases of edge learning that we couldn’t explore in this section, from Keyword spotting [335, 151], spam detection [40, 41], IoT threats prediction [39], Camera Trap Images classification [336], Detecting Defects in photovoltaic Components [337], Estimating Air Quality [338], to Face Spoof Attack Detection [339] and speech recognition [340, 90, 91, 341, 342, 134, 89], another interesting potential application explored in [343] is the usage of Edge Learning in Lunar analogue environments for future space missions. In general, any use case that benefits from personalization on private user data, or suffers from bandwidth limitations or privacy risks in the training phase might benefit from fully or partially using edge learning. Therefore, we expect the trend of edge learning to continue rising, expend into other fields and areas and grow beyond the current usecases and applications in both academia and the industry.
VIII Libraries, Simulators and Tools for Edge Learning
As an emergent field, Edge Learning (EL) requires multiples tools to facilitate its usability, integration and implementation, ranging from emulators and simulators, used train and test ML models on cloud servers before training on the edge, to libraries that allows the successful training of ML models on edge devices.
VIII-A Libraries and Frameworks for edge learning in edge devices
Although there has been significant work on creating libraries and frameworks for ML at the edge, most of these libraries are primarily focused on the deployment and inference of Deep Learning (DL) on edge devices [344], and only a relative few enable the training of ML models in edge devices. Those tools and libraries include ONNX Runtime [345], TensorFlow Lite [346] or libraries that focus on distributed learning tasks such as flower [347] or FedML [348].
The most popular frameworks for training deep learning models, PyTorch [349] and TensorFlow [350] have both developed libraries for edge learning. However, while TensorFlow lite [346], allow for both the inference and training of DL models at the edge, PyTorch mobile [351] and ExecuTorch [352], edge learning libraries for PyTorch at the edge, only allow for inference at the time of writing this article. However, it is possible to train PyTorch models using the ONNX Runtime [353]. ONNX Runtime is a cross-platform inference and training machine-learning accelerator, with On-Device Training capabilities [345, 354], the tool has deep integration with PyTorch [353], as well as other deep learning libraries such as Hugging Face [355] and TensorFlow to accelerate training and inferencing on multiple platforms including mobile devices (Android, IOS) and supports multiple hardware accelerators and programming languages.
Federated learning is the most adopted technique to train ML models at the edge (see section V-E), over the years multiple tools and libraries have been proposed to train FL algorithms on the edge. TensorFlow has developed TensorFlow Federated [356] an open-source framework for machine learning and other computations on decentralized data, however, TensorFlow Federated can’t be used to train models on edge devices at the moment of writing this article and is currently intended for experimentation and research uses only. FL_PyTorch [357], a suite of open-source software built on top of PyTorch [349] for Federated Learning, is another library that focus on research, prototyping and experimenting but does not include FL Runtime support [357], other similar libraries include FedJax [358], LEAF [359] and FedLab [360].
Libraries that support the experimentation and development of FL Algorithms, as well running, training and the deployement of these algorithms in edge devices, include the afformentioned ONNX Runtime that support FL on edge devices as part of its on-device training support [345, 354]. Flower [347] is a library that is aimed toward research and experimentation of FL algorithms, but allow the execution of the algorithms on a variety of edge devices. FedML [348] is a yet another library that allow researcher to experiment with FL algorithms and migrate them to edge devices. PySyft [361] a FL open-source library built as an extension of some DL libraries, such as PyTorch, Keras and Tensorflow, and allow to be run on mobile devices using KotlinSyft  [^3] KotlinSyft: Syft worker for secure on-device machine learning for Android https://github.com/OpenMined/KotlinSyft for Android and SwiftSyft  [^4] SwiftSyft: Syft worker for secure on-device machine learning for iOS https://github.com/OpenMined/SwiftSyft for iOS. And finally, FedERA [362] is a similar library that also includes a verification Module for ensuring the validation of local models to avoid aggregating malicious local models and a carbon emission tracker module to accurately estimate CO2 emissions during the local parameter update phase.
Edge learning frameworks and libraries that are aimed at specific platforms or hardware have also been proposed. They include CoreML [363], an ML inference and training Framework usable in swift for IOS, proposed by Apple to simplify the deployment and training of ML models on IOS environment. And aimed at MCUs, PULP-TrainLib was proposed in [364] to enable On-Device Training for RISC-V Multi-core MCUs. Librairies and frameworks for other approaches then FL include SplitEasy [101] a framework for training split learning algorithms on mobile devices. And EdgeRL [246] for On-Device Reinforcement Learning, a Light-Weight C/C++ Framework, designed to run on a single-core processor that is typically included in a resource-limited embedded platform.
VIII-B Simulation tools and emulators for edge learning
Emulators and simulation tools are both important in the context of edge computing. simulation tools are used to model the behaviour of fog/edge infrastructures, while emulators provide a way to emulate realistic functionality and configurations, allowing for the study of interoperability across different layers and protocols in edge-cloud environments [365]. In the context of EL, simulators and emulators, allow for the experimentation of ML models on cloud servers or a single device, while emulating the resources, memory and energy constraints of real edge devices enabling rapid prototyping and experimentation.
Most simulation tools available in the context of edge learning are aimed at federated learning, Flower [347], FedML [348], PySyft [361] and FedERA [362] that have been presented in the section VIII-A,allow running FL algorithms on edge devices in addition to the simulation. FL_PyTorch [357] is a PyTorch [349] based simulation tool for federated learning, and TensorFlow Federated [356] is a similar tool for the TensorFlow framework [350]. Others Federated learning simulation tools include : LEAF [359]; FedJax [358] a JAX-based open source library; Flute [366] an open source platform with novel optimization, privacy, and communications strategies ; And finally FedLab [360] a lightweight open-source framework that focus on algorithm effectiveness and communication efficiency, and allows customization on server optimization, client optimization, communication agreement, and communication compression.
IX Open issues, research directions and Future trends
In this section, we dive into the challenges, upcoming research paths, and future trends for Edge Learning (EL). Each part focuses on specific issues, shedding light on the evolving landscape. From handling resource constraints to exploring privacy-friendly applications and combining different techniques, these insights offer a concise look into the forefront of edge learning research.
IX-A Run-time optimization
The resource constraints inherent in edge devices, marked by limited memory, computing capabilities, and often restricted energy availability in mobile edge devices, pose substantial challenges for training ML models. Throughout this paper, we have delved into various techniques and approaches designed to optimize and accelerate the training of ML models on edge devices. However, the focal point of future research around edge learning is anticipated to revolve around the refinement and innovation of these optimization strategies. Given the critical need to mitigate the resource demands of ML training on edge devices, the main research direction is expected to be the development of novel techniques that demand less memory, energy, and computational power. This imperative shift towards more efficient runtime optimization is crucial for ensuring the feasibility and scalability of machine learning on edge devices, ultimately facilitating broader adoption across diverse applications and scenarios.
IX-B Growing Usage of Distributed Methods for Edge Learning
As described in Section V-A, distributed learning methods, such as Federated Learning (FL) and Split Learning (SL), stand out as pivotal techniques in the landscape of edge learning. These methods offer a compelling advantage by distributing the training load of large models across multiple devices, alleviating the computational burden on individual edge devices. Current trends, exemplified in Figure 5, show a steady increase in the adoption of FL and SL in the context of edge learning, with FL emerging as the predominant approach for training ML models at the edge. The ongoing trajectory of this trend is expected to persist, driven by the diverse benefits offered by distributed learning methods. This growing reliance on distributed methodologies not only addresses the challenges posed by resource limitations on edge devices but also paves the way for scalable and efficient ML applications in decentralized environments.
IX-C Hybridization of ML techniques in Edge learning
In the exploration of EL techniques detailed in Section V, various strategies have been employed to optimize the training of relatively large ML models on edge devices. However, each of these techniques, as outlined in Table II, comes with distinct advantages and drawbacks. Recognizing the diversity in these methodologies, there has been a notable surge in approaches that advocate for a hybridization of multiple techniques. Figure 6 illustrates this promising trend, wherein researchers aim to maximize the advantages and mitigate the drawbacks of individual techniques by combining them. This emerging direction, signifies a deliberate effort to create comprehensive and robust solutions tailored to the unique challenges posed by edge devices. Given the promising results showcased by such hybrid approaches, the trajectory indicates a continued surge in interest and research efforts towards refining and expanding the applicability of hybridized techniques in the domain of edge learning.
IX-D Lack of labelled data availability
In the context of training ML models at the edge, a prominent challenge arises from the inherent abundance of unlabeled data gathered by edge devices. This issue becomes particularly pertinent as the majority of ML applications traditionally emphasize supervised learning paradigms, necessitating labeled datasets for effective model training and adaptation on edge devices [32]. To address this challenge and pave the way for future research, it is imperative to explore innovative directions. This includes delving into unsupervised (section VI-A), self-supervised (section VI-D), or semi-supervised (section VI-C) techniques and exploring methodologies that can make the most of limited labeled instances such as few-shot learning. Moreover, the development of auto-labeling systems, exemplified by solutions like Flame [367], emerges as a promising avenue to autonomously generate labeled data, mitigating the impact of the labeled data scarcity at the edge.
IX-E Reducing Energy consumption and Carbon Footprint
The surge in distributed learning methodologies, while propelling advancements in ML, has concurrently raised concerns regarding their environmental impact. The substantial energy requirements for training models and data transfer to/from centralized data centers contribute to a significant carbon footprint [368, 369]. Recent trends in machine learning underscore the critical need to estimate and minimize the environmental impact model training processes. To contextualize the magnitude of this issue, According to [370], the estimated carbon footprint associated with edge devices by 2027 will be between 22 and 562 MtCO2-eq/year. Therefore, as a pressing research direction, there is a growing emphasis on developing techniques for training ML models at the edge with meticulous attention to energy efficiency and carbon footprint reduction. Pioneering works, such as [371], have initiated analyses to quantify the environmental impact of ML in edge devices. Notably, frameworks like FedERA [362], designed for training FL models at the edge, incorporate a dedicated carbon emission tracker module to precisely estimate CO2 emissions during the local parameter update phase. More research and development into this, ensures that advancements in Edge ML align with sustainability goals, making it an essential facet of ongoing research in the field.
IX-F Security concerns for distributed learning methods
The advent of distributed learning methodologies, while transformative, introduces notable security concerns, particularly regarding the vulnerability of edge devices [372, 373] and the inherent challenges in securing IoT and mobile devices further compound these concerns [374]. As the cybersecurity landscape for ML models continues to evolve, with a growing literature on the subject (e.g., see [375]), there is a pressing need for comprehensive research into the security implications of training ML models at the edge. Recent studies, exemplified by works like [376, 377, 372, 373], underscore the urgency of understanding potential security vulnerabilities in EL methodologies. Addressing these concerns necessitates further exploration of security measures to be integrated into the training processes of machine learning at the edge, ensuring a robust and secure deployment of distributed learning methods.
IX-G Training Large models at the edge
Large models, including Large Language Models (LLMs) [378, 379, 380], Diffusion models [381], and Audio Generation models [382, 383], etc, are increasingly prevalent and are steadily growing in popularity. However, despite a growing demand for personalization of these models [384, 385] and privacy concerns in collecting and using personal or private data on centralized cloud servers [386]. The fine-tuning, training or personalization of such models at the edge is very challenging considering the limited resources available for edge devices [387]. Although contributions in this domain are currently limited, the predicted surge in interest prompts a need for proactive exploration. Notable approaches, such as FedLLM  [^5] FedLLM is a platform to Build Large Language Models on Proprietary Data using federated learning using the FedML Platform https://doc.fedml.ai/federate/fedllm, FwdLLM [81] and FATE-LLM [80], have emerged using FL to address the challenges of training LLMs at the edge. Looking ahead, the increasing popularity of LLMs and diffusion models anticipates a growing interest in adapting them for EL. Furthermore, innovative techniques such as LoRa [388] and Fnet [389] offer potential solutions for the resource constraints on edge devices, especially when integrated with complementary approaches like FL V-A1, SL V-A2, or model compression techniques V-C. The convergence of these methodologies holds promise for overcoming challenges associated with training large models at the edge in the foreseeable future.
IX-H Extension to privacy preserving application
The escalating popularity of ML applications has brought forth heightened concerns regarding the vast amounts of private and personal data required for effective model training, raising questions about various legal and ethical implications. As discussed in earlier sections, edge learning emerges as a potential solution to address these privacy concerns, as it enables the training of ML models directly on the edge device, eliminating the need for sensitive data to traverse external networks. As such, the usage of edge learning for privacy-preserving applications is expected to be a pivotal research direction for the field. domains like healthcare (VII-A) often had legal requirements as well as ethical concerns of using the data for training ML model [390]. And as explored in discussed in previous sections (VII-D), recommendation systems also stand out as a promising avenue for exploration because the significant scrutiny faced for their reliance on private data during model training [312, 311]. Additionally, other applications that require model personalization and tuning on private data ranging from spam detection in SMS and emails to word suggestions in keyboards and personal assistant chatbots or HAR, can benefit from edge learning, fostering a paradigm shift towards more ethical and privacy-conscious ML applications.
IX-I Frameworks to implement training
Despite the proliferation of frameworks and libraries aimed at enabling ML on edge devices, the current landscape lacks robust support for on-device training. Existing tools, such as PyTorch Mobile [351] and ExecuTorch [352], predominantly emphasize inference on mobile/edge devices, neglecting the essential backpropagation algorithms crucial for the training phase. This imbalance in focus between training and inference highlights a critical gap in the current ecosystem. Although some tools have been proposed to facilitate on-device training at the edge VIII-A, there is a pressing need for the development of new libraries, frameworks, and tools explicitly designed for edge learning. Foreseeing the imminent surge in interest, we anticipate an influx of innovative solutions in the coming years, addressing the scarcity of options for on-device ML training. Furthermore, this surge in interest is likely to catalyse advancements in Internet of Things (IoT) and mobile hardware, optimizing their design and capabilities for ML training on the edge.
X Conclusion
In this survey, we tried to cover the extensive field of Edge learning (EL), which we define as the training and tuning of ML models in the edge, we made an attempt to define the metrics and requirements of EL, and explored then compared the different techniques and methodologies used to optimize ML training at the edge. Next we explored the usage and integrations of ML types such as unsupervised learning, reinforcement learning, etc, and the different applications and use cases of EL. We then explored the tools, libraries and frameworks used for simulating EL and training ML models in edge devices. Finally we identified key challenges in EL and attempted to predict the future trends and research directions of the field. For the techniques and methodologies used to optimize EL, we identified that distributed learning methods such as Federated Learning (FL) are more popular and still growing in popularity for EL, which comes as no surprise considering the potential of these techniques in terms of training large ML models in resource constrained devices. We also estimated the benefits and drawbacks of most of the techniques presented in the survey and identified the growing trend of combining different techniques to alleviate their negatives and make the most out of their benefits. However, the scope of this survey is on the broad concept of EL rather than providing an in depth evaluation and comparison of the performances on specific tasks and problems, as such, more research on specific domains, tasks and techniques is required in the future to complement this survey, and this paper can be used as a reference for developing a broad understanding of edge learning, its requirements, challenges, uses cases and trends, as well as, identifying and gaining a broad overview of the principal techniques used to optimize ML models for training at the edge and pinpointing the different tools, simulations, libraries and frameworks for the training in edge devices.